OpenStudy (anonymous):

verify that parallelogram ABCD with vertices A (-5 -1),B(-9,6),C(-1,5) and D(3,-2) is a rhombus by showing it is a parallelogram with perpendicular diagonals

OpenStudy (anonymous):

Help please :)

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

i think the vertices are wrong? that is not a rhombus

OpenStudy (anonymous):

the product of the gradient of the diagonals must equal to -1 for them to be perpendicular

OpenStudy (anonymous):

IDK, that's the question though

OpenStudy (anonymous):

crazy huh?

OpenStudy (anonymous):

questions wrong then. :/

OpenStudy (anonymous):

I copied and pasted, don't know how to answer it, can you help at all?

OpenStudy (anonymous):

Anyone??

OpenStudy (anonymous):

Can someone else give it a try????

OpenStudy (anonymous):

Please?????

OpenStudy (anonymous):

A (-5 -1),B(-9,6),C(-1,5) and D(3,-2) midpoint of AC is (-5-1)/2 ,(-1+5)/2 --> (-3,2) midpoint of BD is (-9+3)/2,(6-2)/2 --> (-3,2) thus the diagonals AC and BD bisect each other hence ABCD is a ||gm

OpenStudy (anonymous):

now slope of AC is (5-(-1))/(-1-(-5)) =6/4=3/2 again slope of BD is (-2-6)/(3-(-9)) = -8/12 =-2/3 slope of AC * slope of BD =3/2 *-2/3 =-1 hence AC and BD are at right angles bcoz ABCD is a ||gm and the diagonals bisect each other at right angles it is a rhombus